# AutoTimes: Autoregressive Time Series Forecasters via Large Language Models

### v2

@@ Appendix E Showcases
To facilitate a clear comparison among various models, we present additional prediction showcases for long-term forecasting and short-term forecasting. These examples are provided by the following models: Time-LLM [15], FPT [48] and PatchTST [26]. Of all the models, AutoTimes delivers the most accurate future series predictions. Additionally, we provide the showcases of zero-shot and in-context forecasting in Figure 11.
@@ F.2 Impact on Future Research
In this paper, we find prevalent non-autoregressive LLM4TS methods have inconsistencies in the model structure and generative approach with LLMs, leading to insufficient utilization of the inherent multi-step token transition. Given that the generalizability and generative ability of LLMs are largely derived from the autoregressive manner, the potentials of LLMs may not be fully exhibited in time series forecasting. Therefore, we propose to adapt LLMs by the consistent training objective, the next token prediction, and accomplish arbitrary-length forecasting by iterative generation. Beyond the conventional forecasting paradigm, we propose in-context forecasting, where the context for prediction is extended, and earlier historical time series can be utilized as advantageous prompts. The compatibility with LLMs and insights from autoregression can be engaging for future LLM4TS research and the development of foundation time series models.

### v4

@@ Appendix E Showcases
To facilitate a clear comparison among various models, we present additional prediction showcases for long-term forecasting and short-term forecasting. These examples are provided by the following models: TimeLLM [15], FPT [49] and PatchTST [26]. Of all the models, AutoTimes delivers the most accurate future series predictions. Additionally, we provide the showcases of zero-shot and in-context forecasting in Figure 12.
@@ F.2 Impact on Future Research
In this paper, we find prevalent non-autoregressive LLM4TS methods have inconsistencies in the model structure and generative approach with LLMs, leading to insufficient utilization of the inherent multi-step token transition. Given that the generalizability and generative ability of LLMs are largely derived from the autoregressive manner, the potentials of LLMs may not be fully exhibited in time series forecasting. Therefore, we propose to adapt LLMs by the consistent training objective, the next token prediction, and accomplish arbitrary-length forecasting by iterative generation. Beyond the conventional forecasting paradigm, we propose in-context forecasting, where the context for prediction is extended, and earlier historical time series can be utilized as advantageous prompts. The compatibility with LLMs and insights from autoregression can be instructive for future LLM4TS research and the development of foundation time series models.